# Query of CC: Unearthing Large Scale Domain-Specific Knowledge from Public Corpora

### v3

@@ 4.1 Query Analysis
Figure 3: The category distribution of the query for Knowledge Pile.
The progress of query bootstrapping initiates from some categories. Inspiration of the classification of Hendrycks et al. (2021a), we select multiple categories for our initial seed information in the STEM (Science, technology, engineering, and mathematics), Humanities sciences, Social Sciences, and miscellaneous. The key keywords for each category are as follows:
@@
  • Misc: medicine, virology, commonsense knowledge and other miscellaneous.
@@
4.2 Data Statistics
@@
Table 2: Top 10 web domain of Knowledge Pile
@@
Table 3: Main results of our model and baselines in some mathematical reasoning tasks and knowledge related reasoning tasks.
Based on the proposed method, we have formed a high-quality knowledge dataset Knowledge Pile, which maintains about 735GB disk and 188B tokens (using Llama2 tokenizer). As shown in Figure 1, comparing with other datasets in academic and mathematical reasoning domains, we have acquired a large-scale, high-quality knowledge dataset at a lower cost, without the need for manual intervention. Through automated query bootstrapping, we efficiently capture the information about the seed query. Knowledge Pile not only covers mathematical reasoning data but also encompasses rich knowledge-oriented corpora spanning various fields such as biology, physics, etc., enhancing its comprehensive research and application potential.
Table 2 presents the top 10 web domains with the highest proportion of Knowledge Pile, primarily including academic websites, high-quality forums, and some knowledge domain sites. Table 4 provides a breakdown of the data sources’ timestamps in Knowledge Pile, with statistics conducted on an annual basis. It is evident that a significant portion of Knowledge Pile is sourced from recent years, with a decreasing proportion for earlier timestamps. This trend can be attributed to the exponential growth of internet data and the inherent timeliness introduced by the Knowledge Pile.
Figure 4: Timestamp Statistics of Knowledge Pile
5 Experiment and Discussion
5.1 Implementation Details
Based on Knowledge Pile, we further train two models: Llama2-QoC and Mistral-QoC, based on Llama2 and Mistral.
Training
In our experiment, we employed the InternLM [^1] https://github.com/InternLM/InternLM InternLM-Team (2023) library for training all models on 256 A800 GPUs with bfloat16 mixed precision, and only utilized data parallelism during the training process. To enhance throughput and reduce memory consumption, we introduced the Flash attention 2 Dao (2023) module. Above these, both Llama-QoC and Mistral-QoC achieve 4000 tokens per GPU per second(TGS). More training details will be described in Appendix A.
Evaluation
During the evaluation, we utilized the open-source library OpenCompass  [^2] https://github.com/open-compass/opencompass, which serves as a platform for evaluating large models. Leveraging Opencompass, we compare the performance with some open source pre-trained models: Llama2 Touvron et al. (2023b), Code-Llama Rozière et al. (2023), Baichuan 2-Base Yang et al. (2023), Mistral Jiang et al. (2023), Qwen 2 Bai et al. (2023) and some language models for mathematical reasoning: Llemma Azerbayev et al. (2023) and Minerva Lewkowycz et al. (2022). For the selection of evaluation datasets, we opted for three distinct capabilities to assess both Llama2-QoC and Mistral-QoC. These encompassed mathematical reasoning datasets such as Math, GSM8K, knowledge-oriented language understanding datasets including MMLU, AGIEval, and challenging reasoning tasks BIG-Bench hard. More details for evaluation will be described in App B.
5.2 Main Results
Figure 5: The results in MMLU benchmark
The main results of our two models trained on Knowledge Pile ( Llama2-QoC and Mistral-QoC) and the baseline are compared in Table 3 across several general benchmarks. Overall, both models exhibit significantly improved performance, particularly Mistral-QoC. In the complex mathematical reasoning benchmark MATH dataset, Mistral-QoC demonstrates a notable enhancement after QoC training, rising from 11.22 to 17.48, which surpasses professional models such as LLEMMA and Minerva by 3 points (14.3 vs 17.48). Furthermore, Mistral-QoC achieves even higher performance on the mathematical application problem GSM8K (47.31 vs 55.27). Turning to various knowledge-based reasoning tasks, Mistral-QoC displays outstanding capabilities in both MMLU and AGIEval. On the challenging BIG-Bench Hard evaluation set, the model also exhibits a noteworthy improvement in handling complex reasoning tasks.
In comparison to the backbone model, LLAMA-QoC and Mistral-QoC show substantial improvements in mathematical and knowledge-based reasoning tests. For instance, Mistral achieves a 6-points improvement in the MATH dataset and a 5-point improvement in GSM8K. In knowledge-based tests, MMLU shows a relatively modest improvement, within a 1.7 point. However, in AGIEval, Mistral-QoC outperforms Mistral by an impressive 13 points.
@@ 5.2 Main Results
5.3 Knowledge related reasoning
In order to compare the effects of Knowledge Pile on knowledge-related reasoning abilities, Figure 5 contrasts the improvement based on Llama2 and Mistral on MMLU, with a specific focus on performance across different subjects.
Figure 6: Some results in AGIEval benchmark
@@ 5.3 Knowledge related reasoning
From the perspective of Llama-QoC and Llama, following training with Knowledge Pile, the average performance on MMLU for all categories increased by at least 10 points. This notable performance enhancement can be attributed in part to the initially lower performance of Llama, while the Knowledge Pile exhibits relatively higher quality and rich knowledge content, playing a significant role in driving the improvement of Llama2’s performance.
In comparison to Llama, Mistral’s performance improvement is relatively moderate. The results indicate that in the STEM field, Mistral-QoC demonstrates a higher performance improvement compared to other fields, rising from 53.23 to 56.6. In contrast, performance improvements in other categories hover around one point. Two possible reasons account for this phenomenon. Firstly, Mistral exhibits strong comprehension abilities in disciplines such as humanities and social sciences, consistently scoring above 65, leaving relatively limited room for improvement. In contrast, the model’s score in the STEM field is only 53.23, suggesting a greater potential for performance enhancement. Secondly, Knowledge Pile has the largest and most extensive dataset in the STEM field, contributing to the more pronounced performance improvement of Mistral-QoC on MMLU-STEM.
Figure 6 compares the Performance of professional and academic exams covered in AGIEval. Even in the case of professional academic exams, Knowledge Pile similarly leads to substantial performance improvements. As observed in the figure, apart from a slight decline in the SAT-English expression test (from Mistral 68.93 to 65.53), other tasks exhibit noticeable improvements after continued training in Knowledge Pile. Notably, SAT-Math rises from 31.36 to 45, which is a significant improvement for Mistral. Legal exams such as JEC-QA-CA and JEC-QA-KD also show significant enhancements, with the performance increasing from 18.2 and 15.3 to 26.1 and 15.3. For other law-related exam, LSAT is divided into three aspects: Law-Analytics (LSAT-AR), Law-Logic(LSAT-LR), and Law-Reading(LSAT-RC) improvements of 2.17, 8.82, and 7.07 point, respectively. For the LogiQA, Mistral-QoC achieves an accuracy of approximately 45.01%, nearly a 20% improvement compared to Mistral (39.32).
5.4 Difference between Improve Llama and Mistral
As shown in Table 3 and Figure 5, it is evident that Llama2 exhibits inferior performance relative to Mistral. However, this also highlights a greater potential for performance improvement when training in Knowledge Pile. Also, we observe a significant behavioral difference during the improvement process, which is shown in Figure 7. We found that in certain tasks, Mistral’s performance undergoes a certain degree of decline during the improvement process, followed by a subsequent ascent after some time, eventually surpassing the previous performance levels. This phenomenon may come from a conflict between the potential distribution of the model and the distribution of high-quality datasets. Mistral’s potential distribution is superior but more unstable, whereas Llama’s performance is relatively poorer but exhibits greater plasticity. On the other hand, the improvement achieved on more powerful models (Mistral 7B) also demonstrates the high-quality of Knowledge Pile.
@@ 6 Conclusion
In this study, we propose an efficient method Query of CC, for the automated collection of specialized domain data. Leveraging seed data from some specific domains, we employ a large language model for query bootstrapping. By optimizing the breadth and depth of queries, we expand the query to retrieve data relevant to the specified domain. Ultimately, we collected and released an open dataset comprising approximately 735GB of Knowledge Pile, equivalent to approximately 188 billion tokens, in the fields of mathematics and knowledge. Experimental results demonstrate that the adoption of Knowledge Pile significantly enhances the model’s performance in some reasoning tasks, such as math word problems and professional examinations. Our objective is not only to establish a research foundation for community studies in mathematical and knowledge-related reasoning but also to provide an efficient and cost-effective method for collecting high-quality data, thereby facilitating the accumulation of more high-quality data.
7 Limitation
In this paper, we propose a collection method for domain-specific data, Query of CC, which extends domain-relevant queries through LLM bootstrapping for data retrieval. Additionally, we demonstrate through empirical evaluation on further training of LLAMA and Mistral that data collected using this method significantly improves the ability of LLM in some specific domains. However, this method still has the following potential limitations:
Data Quality
The quality of data collected by Query of CC depends largely on the quality of data in public corpora. In this work, we utilized the Common Crawl corpus as our public corpus, extracting and processing the CC dump up to April 2023. Despite leveraging some methods Wenzek et al. (2020) for processing high-quality web data, we found that the corpus still contains an amount of low-quality and erroneously extracted content. Thus, improving the data quality of public corpora is also a direction for future research.
Data Scale
The data source of Query of CC is limited by public corpora, and unable to collect information not present in these corpora. This limitation requires particular attention. Moreover, the proportion of domain data in public databases also influences the scale of the collected dataset. During the collection process, scoring based on similarity scores can be employed to filter highly relevant data.

### v4

@@ 4.1 Query Analysis
Figure 3: The category distribution of the query for $\mathsf{Retrieve}$-$\mathsf{Pile}$.
The progress of query expanding initiates from some categories. Inspiration of the classification of Hendrycks et al. (2021a), we select multiple categories for our initial seed information in the STEM (Science, technology, engineering, and mathematics), Humanities sciences, Social Sciences, and miscellaneous. The key keywords for each category are as follows:
@@
  • Misc: medicine, virology, commonsense knowledge.
@@
4.2 Data Analysis
Overview
Based on Retrieve-from-CC, we have formed a high-quality knowledge dataset $\mathsf{Retrieve}$-$\mathsf{Pile}$, which maintains about 735GB disk and 188B tokens (using Llama2 tokenizer). As shown in Figure 1, comparing with other datasets in academic and mathematical reasoning domains, we have acquired a large-scale, knowledge-related dataset at a lower cost, without the need for manual intervention. Through automated query expanding, we efficiently capture the information about the seed query. $\mathsf{Retrieve}$-$\mathsf{Pile}$ not only covers mathematical reasoning data but also encompasses rich knowledge-oriented corpora spanning various fields such as biology, physics, etc., enhancing its comprehensive research and application potential.
Web Domain composition of $\mathsf{Retrieve}$-$\mathsf{Pile}$
@@ 4.2 Data Analysis
Table 2: Top 10 most web domain of the data in $\mathsf{Retrieve}$-$\mathsf{Pile}$, most of these are academic institutions, high-value forums, and authoritative website.
Figure 4: Left: The frequency distribution of the documents number across URL domains, with most domains having few documents, while a small number have many. The y-axis uses a logarithmic scale to highlight this imbalance. this means Retrieve-from-CC not only retrieve the data from high knowledge density websites like Wikipedia but collect data from scatted websites. Right: The timestamp statistics of $\mathsf{Retrieve}$-$\mathsf{Pile}$, most data of $\mathsf{Retrieve}$-$\mathsf{Pile}$  come from recent years (different colors represent different years).
Table  2 presents the top 10 web domains with the highest proportion in $\mathsf{Retrieve}$-$\mathsf{Pile}$, which cover a wide range of academic institutions, high-value forums, and authoritative websites in specific knowledge fields. These resources are closely related to the knowledge domains we aim to collect, such as en.wikipedia.org and www.semanticscholar.org. Many previous research (Touvron et al., 2023a; Gao et al., 2021; Taylor et al., 2022) have specifically collected data from these domains to enrich the knowledge of the training dataset. To gain an insight into the data distribution of $\mathsf{Retrieve}$-$\mathsf{Pile}$, we randomly selected 100,000 examples and conducted the statistical analysis of their domain frequency. Figure 4 left shows the frequency distribution of the document number across the URL domain. We observed that in the $\mathsf{Retrieve}$-$\mathsf{Pile}$, the vast majority of web domains are recorded only once, and these domains also contain rich knowledge content. However, traditional manual data collection methods have limitations in systematically collecting these scattered data, and Retrieve-from-CC has shown its excellent data collection capabilities in this regard.
Furthermore, Table 4 right statistic the timestamps of data sources in $\mathsf{Retrieve}$-$\mathsf{Pile}$ by year. It is evident that most of the data in $\mathsf{Retrieve}$-$\mathsf{Pile}$  originates from recent years, and the proportion of earlier timestamps is gradually decreasing. This phenomenon can be attributed to the exponential growth of internet data volume and the inherent timeliness characteristic of the $\mathsf{Retrieve}$-$\mathsf{Pile}$.
Data Quality Analysis
Data quality is a highly complex concept. Previous studies (Brown et al., 2020; Jiang et al., 2023; InternLM-Team, 2023; Touvron et al., 2023a) labeled certain portions of their datasets as high-quality and employed trained scorers to quantify dataset quality. In our assessment of the data quality in $\mathsf{Retrieve}$-$\mathsf{Pile}$, we leverage utilized (Wettig et al., 2024), which rates the data quality across four dimensions: writing style, required expertise, facts & trivia, and educational value. Also, a 1.3-billion parameter language model was trained using a pair-wise method to assess the quality of the data. Figure 5 illustrates the distribution of QuRating for the $\mathsf{Retrieve}$-$\mathsf{Pile}$ and The Pile (Gao et al., 2021), alongside representative subsets of The Pile, including Wikipedia, which scores high on factual, Books3, which exhibits a wide variety of writing styles, and Arxiv, which requires a high level of expertise. As can be seen in the figure, $\mathsf{Retrieve}$-$\mathsf{Pile}$ receives high ratings across all dimensions compared to the full Pile dataset. Moreover, in terms of educational value, $\mathsf{Retrieve}$-$\mathsf{Pile}$ demonstrates the highest score.
In addition, we employed an open-source data quality classifier [^2] https://huggingface.co/kenhktsui/llm-data-textbook-quality-fasttext-classifier-v2, to rate the data within $\mathsf{Retrieve}$-$\mathsf{Pile}$. Inspired by Gunasekar et al. (2023), high-quality data should possess characteristics of high educational value, namely: clarity, independence, instruction, and balance. To achieve the assessment of the educational value of the data, Tsui (2024) collected a subset of high-quality raw data and trained a classifier for evaluating the educational value of data based on fasttext  [^3] https://fasttext.cc/. The educational value ranges from 0, indicating low educational value, to 2, indicating high educational value. In Table 3, we present a comparison between $\mathsf{Retrieve}$-$\mathsf{Pile}$ and other mainstream knowledge datasets. The results show that $\mathsf{Retrieve}$-$\mathsf{Pile}$ has an average score of 1.29, significantly outperforming other open-source knowledge-based real datasets.
Figure 5: The distribution of QuRating (Wettig et al., 2024) of $\mathsf{Retrieve}$-$\mathsf{Pile}$, The Pile, and selected high-quality subsets of The Pile. QuRating is a robust metric designed to evaluate data quality across four dimensions, with higher scores indicating better quality. Following Wettig et al. (2024), the scores are normalized to have a mean of zero and a standard deviation of one for all displayed data.
Figure 6: The distribution of educational value of different open source datasets, which shows that the distribution of $\mathsf{Retrieve}$-$\mathsf{Pile}$on the x-axis is significantly right shifted compare others, indicating that it has higher educational value.
Datasets Educational Value ($\uparrow$)
MATHPILE (Wang et al., 2023c) 1.25
Guanaco* (Dettmers, 2023) 1.115
OpenWebMath (Paster et al., 2024) 1.089
Proof of Pile Azerbayev et al. (2024) 1.13
The Pile (Gao et al., 2021) 1.011
RedPajama* (Computer, 2023) 0.985
$\mathsf{Retrieve}$-$\mathsf{Pile}$ (ours) 1.291
Table 3: The comparison of average educational value scores among different open-source datasets. * denotes the results cited from Tsui (2024), which selected the first 100,000 samples of the dataset, and others randomly selected 100,000 samples.
Figure 6 further reveals the differences in the distribution of educational value among the $\mathsf{Retrieve}$-$\mathsf{Pile}$, Pile, and OpenWebMath datasets. Through comparison, we can observe a distinct rightward shift in the distribution of $\mathsf{Retrieve}$-$\mathsf{Pile}$, indicating that $\mathsf{Retrieve}$-$\mathsf{Pile}$contains a greater amount of data with high educational value, while the proportion of low-value data is relatively lower.
4.3 The Improvement in Knowledge-Related Reasoning Benchmark
For evaluate the improvement of  $\mathsf{Retrieve}$-$\mathsf{Pile}$, we conducted two experiments. Firstly, we further trained two models: Llama2-QoC and Mistral-QoC, both of which are based on Llama2 Touvron et al. (2023b) and Mistral Jiang et al. (2023), each with 7 billion parameters. Additionally, we trained a 1.8 billion parameter language model using the Llama architecture from scratch.
Implementation
In our experiments, we employed the InternLM [^4] https://github.com/InternLM/InternLM (InternLM-Team, 2023) library for training all models on 256 A800 GPUs with bfloat16 mixed precision, and only utilized data parallelism during the training process. To enhance throughput and reduce memory consumption, we introduced the Flash attention 2 (Dao, 2024) module. More training details will be described in the Appendix.
During the evaluation, we utilized the open-source library OpenCompass  [^5] https://github.com/open-compass/opencompass, which serves as a platform for evaluating LLMs. Leveraging Opencompass, we compare the performance with some open source pre-trained models: Llama2 (Touvron et al., 2023b), Code-Llama (Rozière et al., 2023), Baichuan 2-Base (Yang et al., 2023), Mistral (Jiang et al., 2023), Qwen 2 (Bai et al., 2023) and some language models for mathematical reasoning: Llemma (Azerbayev et al., 2024) and Minerva (Lewkowycz et al., 2022). For the selection of evaluation datasets, we opted for three distinct capabilities to assess both Llama2-QoC and Mistral-QoC. These encompassed mathematical reasoning datasets such as Math (Hendrycks et al., 2021b), GSM8K (Cobbe et al., 2021), knowledge-oriented language understanding datasets including MMLU (Hendrycks et al., 2021a), AGIEval (Zhong et al., 2024), and challenging reasoning tasks BIG-Bench hard (Suzgun et al., 2023). All metrics of these benchmark reported in this paper is accuracy. More details for evaluation will be described in the Appendix.
The improvement in further training
@@ 4.3 The Improvement in Knowledge-Related Reasoning Benchmark
Table 4: The performance of our further trained model (Llama 2-QoC and Mistral-QoC) and baselines in some mathematical reasoning tasks and knowledge related reasoning tasks. In this table, all metric is accuracy.
The results of our two models trained on $\mathsf{Retrieve}$-$\mathsf{Pile}$ (Llama2-QoC and Mistral-QoC) and the baseline are compared in Table 4 across several general benchmarks. Overall, both models exhibit significantly improved performance, particularly Mistral-QoC. In the complex mathematical reasoning benchmark MATH dataset, Mistral-QoC demonstrates a notable enhancement after QoC training, rising from 11.22 to 17.48, which surpasses professional models such as LLEMMA and Minerva by 3 points (14.3 vs 17.48). Furthermore, Mistral-QoC achieves even higher performance on the mathematical application problem GSM8K (47.31 vs 55.27). Turning to various knowledge-based reasoning tasks, Mistral-QoC displays outstanding capabilities in both MMLU and AGIEval. On the challenging BIG-Bench Hard evaluation set, the model also exhibits a noteworthy improvement in handling complex reasoning tasks.
In comparison to the backbone model, LLAMA-QoC and Mistral-QoC show substantial improvements in mathematical and knowledge-based reasoning tests. For instance, Mistral achieves a 6-point improvement in the MATH dataset and a 5-point improvement in GSM8K. In knowledge-based tests, MMLU shows a relatively modest improvement, within a 1.7 point. However, in AGIEval, Mistral-QoC outperforms Mistral by an impressive 13 points.
@@
Difference between Improve Llama and Mistral
As shown in Table 4, it is evident that Llama2 exhibits inferior performance relative to Mistral. However, this also highlights a greater potential for performance improvement when training in $\mathsf{Retrieve}$-$\mathsf{Pile}$. Also, we observe a significant behavioral difference during the improvement process, which is shown in Figure 7. We found that in certain tasks, Mistral’s performance undergoes a certain degree of decline during the improvement process, followed by a subsequent ascent after some time, eventually surpassing the previous performance levels. This phenomenon may come from a conflict between the potential distribution of the model and the distribution of high-quality datasets. Mistral’s potential distribution is superior but more unstable, whereas Llama’s performance is relatively poorer but exhibits greater plasticity. On the other hand, the improvement achieved on more powerful models (Mistral 7B) also demonstrates the high-quality of $\mathsf{Retrieve}$-$\mathsf{Pile}$.
@@
MMLU
C4 28.52
$\mathsf{Retrieve}$-$\mathsf{Pile}$ 33.13
Table 5: The comparison of performace which pre-training in $\mathsf{Retrieve}$-$\mathsf{Pile}$ and C4.
Overlap Ratio
MMLU 1.5%
GSM8k 0.0%
MATH 0.013%
BBH 0.0%
Table 6: The data contamination between $\mathsf{Retrieve}$-$\mathsf{Pile}$ and downstream task.
Comparation with other pre-training data
Similarly, we evaluated the impact of $\mathsf{Retrieve}$-$\mathsf{Pile}$ on model performance during the pre-training phase. Table 5 compares the performance of a 1.8B-parameter language model pre-trained on 200B tokens from $\mathsf{Retrieve}$-$\mathsf{Pile}$ with that of C4 (Raffel et al., 2020) (an open-source pre-training dataset curated and filtered from Common Crawl) on MMLU. The 200B tokens for C4 were randomly sampled. The model trained on C4 achieved only 28.52% accuracy on MMLU, whereas the model trained on $\mathsf{Retrieve}$-$\mathsf{Pile}$ showed a 5% point improvement (33.13%) in MMLU compared to C4, which highlights the advantage of $\mathsf{Retrieve}$-$\mathsf{Pile}$over randomly sampled data.
Data contamination
Data contamination analysis is equally important in the study of pre-training corpora. Many research (Chowdhery et al., 2023; Touvron et al., 2023b; Jiang et al., 2024) have defined data contamination as the n-gram overlap between the pre-training corpus and the test set. For instance, PaLM Chowdhery et al. (2023) defines contamination as a 70% 8-gram overlap, while Llama2 Touvron et al. (2023b) considers it contamination if more than 10 tokens overlap between the training and test sets. Inspired by previous works, we utilized the Overlapy codebase to calculate the 8-gram overlap between $\mathsf{Retrieve}$-$\mathsf{Pile}$ and the test set. Table 6 shows the proportion of documents with 8-gram overlaps between $\mathsf{Retrieve}$-$\mathsf{Pile}$ and downstream tasks. It is important to note that this metric calculates direct overlaps, potentially leading to false positives, as overlaps may occur between semantically different contents. As shown in table 6, the overlap across most downstream tasks is relatively minimal, generally below 0.1%. For example, gsm8k and BBH exhibit an overlap rate of 0%, while MMLU, containing a significant amount of conceptual content, shows only about a 1.5% overlap. One potential explanation is that MMLU incorporates substantial knowledge-related questions, particularly general knowledge, which may be reflected by similar descriptions in the pre-training corpus. Additionally, domain-specific knowledge typically consists of specialized terms and standardized expressions, which are prone to repetition. Overall, the low n-gram overlap of $\mathsf{Retrieve}$-$\mathsf{Pile}$ across downstream tasks suggests that contamination in $\mathsf{Retrieve}$-$\mathsf{Pile}$ may not be significant. This also highlights the dataset’s broad adaptability in handling diverse tasks.
5 Limitation and Hallucination Analysis
5.1 Limitation
In this paper, we propose a collection method for domain-specific data, Retrieve-from-CC, which extends domain-relevant queries through LLMs generated for data retrieval. Additionally, we demonstrate through empirical evaluation on further training of LLAMA and Mistral that data collected using this method significantly improves the ability of LLM in some specific domains. However, this method still has the following potential limitations:
Data Quality
The quality of data collected by Retrieve-from-CC depends largely on the quality of data in public corpora. In this work, we utilized the Common Crawl corpus as our public corpus, extracting and processing the CC dump up to April 2023. Despite leveraging some methods (Wenzek et al., 2020) for processing high-quality web data, we found that the corpus still contains an amount of low-quality and erroneously extracted content. Thus, improving the data quality of public corpora is also a direction for future research.
5.2 Hallucination Analysis
The output of LLMs generally exhibits significant hallucination issues. Previous work has shown that LLMs are prone to hallucinations, and using their output directly in training without filtering can exacerbate the model’s hallucinations for certain issues. However, in the data collection pipeline of Retrieve-from-CC, the model is only used to generate queries and is not directly applied in training. Therefore, even if the synthesized queries from the LLMs contain incorrect information, the information retrieved from the corpus based on these incorrect queries is correct. Hallucinatory queries do not lead to the retrieval of incorrect information.
@@ 6 Conclusion
In this study, we propose an efficient method Retrieve-from-CC, for the automated collection of specialized domain data. Leveraging seed data from some specific domains, we employ a language model for query expanding. By optimizing the breadth and depth of queries, we expand the query to retrieve data relevant to the specified domain. Ultimately, we collected and released an open dataset comprising approximately 735GB of $\mathsf{Retrieve}$-$\mathsf{Pile}$, equivalent to approximately 188 billion tokens, in the fields of mathematics and knowledge. Experimental results demonstrate that the adoption of $\mathsf{Retrieve}$-$\mathsf{Pile}$ significantly enhances the model’s performance in some reasoning tasks, such as math word problems and professional examinations. Our objective is not only to establish a research foundation for community studies in mathematical and knowledge-related reasoning but also to provide an efficient and cost-effective method for collecting high-quality data, thereby facilitating the accumulation of more high-quality data.